Now this is really a challenging question and I am lost.
At an annual flower show, 7 different entries are to be arranged in a row.

a) how many different arrangements of the entries are possible?
b) If the owners of the 1st, 2nd, and 3rd place entries will be awarded prizes of $100, $50, and $25 respectively, how many ways can the prizes be awarded?

Answer by stanbon(48558)

At an annual flower show, 7 different entries are to be arranged in a row.
a) how many different arrangements of the entries are possible?
Ans: 7! = 7*6*5*4*3*2*1 = 5040
--------------------------------------------
b) If the owners of the 1st, 2nd, and 3rd place entries will be awarded prizes of $100, $50, and $25 respectively, how many ways can the prizes be awarded?
Ans: 7P3 = 7!/(7-3)! = 7*6*5 = 210
=======================================
Cheers,
Stan H.